# EoRA: Fine-tuning-free Compensation for Compressed LLM with Eigenspace Low-Rank Approximation

### v4

@@ 4.1 Experiments Details
We implement EoRA in PyTorch (Paszke et al., 2017), utilizing the Hugging Face Transformers and Datasets framework (Wolf et al., 2019). All experiments are conducted on a single NVIDIA H100 GPU. We primarily focus on evaluating EoRA for compensating LLaMA2-7B/13B and LLaMA3-8B models, compressed using SparseGPT (Frantar and Alistarh, 2023), a widely adopted pruning method, and GPTQ (Frantar et al., 2023) for quantization. Channel-wise asymmetric quantization is applied across all experiments, and we follow the settings from (Huang et al., 2024) to construct the calibration dataset for both SparseGPT and GPTQ.
@@
We compare EoRA with ZeroQuant-V2 Yao et al. (2024) which proposes using simple SVD for optimizing Eq. 2. Although Activation-aware Singular Value Decomposition (ASVD) Yuan et al. (2023) is designed to replace the entire model with its low-rank decomposition rather than approximating the compression errors, its strategy of incorporating activation distribution variance can also be adapted for error compensation using low-rank matrices. Specifically, we scale the compression error $\Delta W$ using a diagonal scaling matrix $S$, where each diagonal entry $S_{ii}$ is computed based on the average absolute value of the activations $\tilde{X}$ in the $i$-th channel as $S_{ii}=\left(\frac{1}{n}\sum_{j=1}^{n}|\tilde{X}_{ij}|\right)^{\frac{1}{2}}$. Here, $n$ denotes the number of activation entries in the $i$-th channel. We then apply SVD to the scaled error $\Delta W^{"}=\Delta WS$ to obtain its low-rank approximation. Since $S$ is invertible, we can project the approximation back to the original space as $\Delta W=\Delta W^{"}S^{-1}\approx B^{"}A^{"}S^{-1}$—same as how EoRA project its compensation back to the original space. We refer to this method as Act-S in the remainder of this paper. We also compare EoRA with a training-based method, ApiQ Liao et al. (2024), which optimizes low-rank matrices ($A$ and $B$) using gradient-based training to minimize Eq. 6. In our comparison, we limit the evaluation to the layer-wise variant of ApiQ, as other variants require substantially more memory or training time. These more resource-intensive versions align more closely with PEFT methods rather than training-free low-rank approximation approaches. For instance, when applied at the model level, ApiQ effectively becomes equivalent to training LoRA on top of a compressed model—shifting its focus toward fine-tuning rather than training-free compensation, and thus falling outside the scope of this study. Note that the optimization time for both EoRA and Act-S is comparable, with each completing within minutes, whereas ApiQ requires over hours to optimize.
@@
We evaluate EoRA and the baselines on improving the task-specific accuracy of the compressed LLMs on language generation, commonsense reasoning, and math reasoning tasks using the LM-Evaluation-Harness framework (Gao et al., 2024). We pick WikiText2 for the language generation task and perplexity as the evaluation metric. For commonsense reasoning, we select ARC-Challenge (ARC-C) (Clark et al., 2018), and for math reasoning ability, we choose MathQA (Amini et al., 2019) and GSM8K (Cobbe et al., 2021). We sample 128 concatenated sentences of length 2048 from the WikiText2 training set as the calibration set for EoRA, Act-S, and ApiQ for the language generation task. For commonsense reasoning tasks, we sample 32 concatenated sentences of length 2048 from the ARC training set and combine them with 32 concatenated sentences of the same length from C4 (Raffel et al., 2020) to construct the calibration set for EoRA, Act-S, and ApiQ. Similarly, for the math reasoning tasks, we sample 32 concatenated sentences of length 2048 from the MathQA/GSM8K training set and combine them with 32 concatenated sentences from C4 to form the calibration set for the three methods.
@@ 4.2.1 Sparsity Error Compensation
To assess the effectiveness of EoRA in compensating for sparsity error, we compare EoRA with all the baselines on LLaMA2-7B/13B and LLaMA3-8B models pruned with SparseGPT to 2:4 sparsity—the only sparsity format that yields actual inference speedups on GPUs. Rank of all the compensation methods is set to 128, and the results of LLaMA3-8B are presented in Table 1, while the full results, including LLaMA2-7B/13B, are provided in Table 4 in the appendix. EoRA consistently outperforms all training-free baselines, achieving gains of $2.9\%$, $2.1\%$, and $10.7\%$ over Act-S on ARC-C, MathQA, and GSM8K, respectively. Furthermore, it surpasses ApiQ by $0.4\%$ on ARC-C and $1.1\%$ on MathQA, while delivering comparable results on GSM8K—all with significantly faster optimization time (15 minutes vs. 2.5 hours). Furthermore, EoRA proves robustness across different model sizes, continuing to outperform ZeroQuant-V2, Act-S and ApiQ in boosting the accuracy of 2:4 pruned LLaMA2-7B/13B across ARC-C and MathQA as shown in Table 4. We further assess the generalizability and compatibility of EoRA with pruning methods beyond SparseGPT. Specifically, we evaluate EoRA on LLaMA3-8B pruned to 2:4 sparsity using Wanda Sun et al. (2024), where EoRA continues to outperform all the training-free baseline methods. For additional details, please refer to Section A.4.
@@ 4.2.2 Quantization Error Compensation
We evaluate EoRA on LLaMA2-7B/13B and LLaMA3-8B models quantized with GPTQ to 4-bit and 3-bit to assess the effectiveness of EoRA in compensating for quantization error. The ranks for all the methods are set to 128. From Table 2, 3-bit quantization causes significant accuracy degradation, with losses up to $29.5\%$/$17.7\%$/$35.8\%$ on ARC-C, MathQA, and GSM8K, respectively. By applying EoRA, we demonstrate that the accuracy loss can be reduced to $18.7\%$/$10.9\%$/$24.3\%$ on ARC-C, MathQA, and GSM8K—providing $10.8\%$/$6.7\%$/$11.5\%$ improvement, outperforming all the baseline methods for compensating the quantization error. On the other hand, although 4-bit quantization does not result in as much accuracy loss as 3-bit quantization, applying EoRA can still generally enhance the performance of the 4-bit model, offering up to a $2.2\%$ and $3.14\%$ accuracy boost on ARC-C and MathQA, respectively. Comprehensive results, including those for LLaMA2-7B/13B, are presented in Table 5 in the appendix, where a similar trend of improvement with EoRA is observed. We also explore the feasibility of using EoRA to improve ultra-compressed models that combine both pruning and quantization. In this setting, EoRA continues to outperform all baselines on ARC-C and MathQA. Further details can be found in Section 6 of the Appendix.
@@ 4.3 Ablation Study: Ranks and Calibration Sizes
Since one of the advantages of using EoRA is the greater flexibility in adjusting overall model accuracy without being constrained by specific compression formats, in this section, we investigate the influence of different ranks on adopting EoRA. We vary the rank of EoRA in {64,128,256,512} on compensating LLaMA3-8B pruned to 2:4 sparsity. As shown in Figure 2, EoRA consistently outperforms the two training-free baselines (ZeroQuant-V2 and Act-S) across all tested ranks, with the performance gap becoming more prominent at higher ranks. For instance, on GSM8K, EoRA achieves improvements of $7.43\%$, $10.69\%$, $11.9\%$, and $14.62\%$ at ranks 64, 128, 256, and 512, respectively. In contrast, the gains on ARC-C remain relatively stable across ranks, ranging between $2\%$ and $4\%$. Additionally, EoRA begins to outperform ApiQ on GSM8K at higher ranks, with improvements of $1.21\%$ and $2.51\%$ observed at ranks 256 and 512, respectively. These experiments prove that EoRA is robust across different rank settings, offering users a more flexible option upon existing compression configurations to effectively balance the trade-off between inference overhead and model accuracy. A similar trend is observed in the results for LLaMA2-7B/13B shown in Table 8 in the appendix.
We also compare the influence of different calibration sizes on EoRA . We vary the calibration size in {16,32,64}, and compare them on recovering the accuracy of LLaMA3-8B quantized to 3/4-bit and pruned to 2:4 sparsity. Overall, we find that EoRA demonstrates strong robustness and maintains competitive accuracy even with limited calibration data, as shown in Table 9. Notably, using as few as 32 calibration samples to compensate a 2:4 pruned model can even yield better accuracy improvements than using 64 samples.
@@ 4.4 EoRA as LoRA initialization for Fine-tuning Compressed Models
LLaMA3-8B Full-precision - w/o finetuning 50.42 40.10
@@
EoRA 55.46 56.04
We show that, with additional computational resources, users can leverage the low-rank matrices from EoRA as initialization for LoRA fine-tuning, enabling further accuracy improvements for compressed models. We follow the conventional LoRA fine-tuning framework, which keeps the compressed model frozen and only tunes the low-rank residual components during fine-tuning. We conduct experiments on compressed LLaMA3-8B models with {2:4 sparsity, 4-bit, 3-bit} compression. The rank of LoRA is set to 128 and is applied to every linear layer, initialized using EoRA, SVD following LoftQ Li et al. (2024), and standard initialization following QLoRA Dettmers et al. (2023). Fine-tuning is performed on the ARC training set for evaluating ARC-C, and on the MathQA training set for the math reasoning task. We fine-tune the models for 3 epochs with a batch size of 64, a learning rate of 1e-5, and a cosine learning rate scheduler. As shown in Table 3, initializing with EoRA substantially enhances the accuracy of compressed models, surpassing both QLoRA and LoftQ when fine-tuning 4-bit quantized LLaMA3-8B, and achieving accuracy on par with standard full-precision fine-tuning. We also observed that the improvements over QLoRA and LoftQ are more pronounced on 3-bit quantized and 2:4 pruned models, aligning with our earlier finding that EoRA is more effective when the compression error is more substantial, as shown in Table 10.
@@ 4.5 Kernel Optimization, Inference Speed Evaluation and Memory Overhead of EoRA
While theoretically, compensating a compressed model with low-rank residual paths introduces minimal computational overhead, in practice, it leads to a noticeable increase in latency. This is primarily because input and output must transfer between L2 cache and DRAM twice as often compared to that without a low-rank residual path, shifting the inference process from being computation-bound to memory-bound. This phenomenon is also discussed in  (Li et al., 2025). To address this, we propose fusing the low-bit weight quantization kernel with the matrix multiplication of $B$, which shares the same output. By doing so, the shared output no longer needs to be offloaded and reloaded to the L2 cache, effectively reducing data transfer overhead as illustrated in Figure 3 (a). Implementation details of our kernel can be found in Section A.8. As shown in Table 11, our custom EoRA kernel substantially accelerates inference compared to using native PyTorch for the low-rank residual path on top of the low-bit quantized kernel, achieving a speedup of up to 1.4x over FP16 with EoRA of rank 128 at 3-bit quantization. In contrast, without the EoRA kernel, the initial 1.7x speedup provided by the 3-bit quantized kernel drops to 1.1x. Similarly, under 4-bit quantization, the EoRA kernel delivers an extra 0.3x speedup compared to setups without the EoRA kernel.
Finally, EoRA can also be quantized to further reduce the additional cost of residual low-rank compensation paths. In this section, we quantize EoRA of rank {128, 512} to 4/3-bit on compensating three types of compressed LLaMA3-8B models (2:4 pruned, 4-bit quantized, and 3-bit quantized). The complete results are provided in Table 12 in appendix, while the results for LLaMA3-8B are illustrated in Figure 3 (b). As shown in the figure, EoRA is robust to quantization, which means that when EoRA is quantized, the accuracy drop from full-precision EoRA is insignificant while the model size is significantly reduced. For example, when a 512-rank EoRA is quantized from 16-bits to 4-bit on 2:4 pruned LLaMA3-8B, the accuracy drops are only $0.43\%$ on ARC-C while the total model size reduces by $16.49\%$. Additionally, compared to the original uncompensated 2:4 pruned model, quantizing EoRA of rank 128/512 improves accuracy by $4.4\%$/$11.4\%$ with a total model size increase of just $2\%$/$7\%$. For 3-bit quantized LLaMA3-8B compensated with a 4-bit quantized EoRA of rank 128/512 achieves $10.6\%$/$19.1\%$ accuracy improvements, with a corresponding model size increase of only $3\%$/$14\%$. Interestingly, we also observe that quantizing EoRA does not always result in accuracy loss; in some cases, it even slightly improves accuracy, potentially due to quantization acting as a form of regularization, as discussed in Liu et al. (2023). Generally, we recommend users quantize EoRA to 4-bit, as this significantly reduces inference latency and model size with kernel support, without causing any noticeable drop in accuracy.
@@ 5 Related Works
Post-training LLM Compression: As LLMs scale, reducing their size is essential for efficient deployment. Traditional compression-aware training methods are impractical due to the need for full datasets and heavy retraining. Post-training compression methods like quantization and pruning have gained popularity as they require only minimal calibration data and no retraining. PTQ reduces model size by lowering bitwidths (Frantar et al., 2023; Tseng et al., 2024), while PTP removes less important weights to reduce computation (Frantar and Alistarh, 2023; Sun et al., 2024). Our method, EoRA, is compatible with all such compression techniques as it operates independently of the base method used.
Low-Rank Decomposition: Low-rank decomposition methods (Yuan et al., 2023; Sakr and Khailany, 2024; Wang et al., 2025; Hsu et al., 2022) compress models by replacing weights with low-rank matrices, reducing both latency and size without special kernel support. However, they are less widely adopted due to weaker accuracy-compression trade-offs. SVD-LLM (Wang et al., 2025) is conceptually close to our work which also tries to align the SVD compression error with the layer-wise compression loss—it relies on the matrix product of the activation being positive-definite, a condition often unmet in practice. Enforcing this condition typically requires additional modifications, which introduce noise into the approximation. In contrast, EoRA employs eigendecomposition, which only requires the matrix product of the activation to be symmetric—a property that naturally holds—avoiding such issues. Furthermore, although EoRA also utilizes SVD-based low-rank decomposition, its core objective is fundamentally different. Whereas prior methods aim to replace pre-trained weight matrices with low-rank approximations to reduce model size and inference cost, EoRA focuses on approximating the compression error itself. This allows for improved accuracy recovery in compressed LLMs and provides greater flexibility in balancing accuracy and computational overhead by overcoming the constraints of fixed compression formats.

### v5

@@ 4.1 Experiments Details
We implement EoRA in PyTorch (paszke2017automatic), utilizing the Hugging Face Transformers and Datasets framework (wolf2019huggingface). All experiments are conducted on a single NVIDIA H100 GPU. We primarily focus on evaluating EoRA for compensating LLaMA2-7B/13B and LLaMA3-8B models, compressed using SparseGPT (frantar2023sparsegpt), a widely adopted pruning method, and GPTQ (frantar2022gptq) for quantization. Channel-wise asymmetric quantization is applied across all experiments, and we follow the settings from (huang2024good) to construct the calibration dataset for both SparseGPT and GPTQ.
@@
We compare EoRA with ZeroQuant-V2 (yao2024zeroquantv2) which proposes using simple SVD for optimizing Eq. 2. Although Activation-aware Singular Value Decomposition (ASVD) (yuan2023asvd) is designed to replace the entire model with its low-rank decomposition rather than approximating the compression errors, its strategy of incorporating activation distribution variance can also be adapted for error compensation using low-rank matrices. Specifically, we scale the compression error $\Delta W$ using a diagonal scaling matrix $S$, where each diagonal entry $S_{ii}$ is computed based on the average absolute value of the activations $\tilde{X}$ in the $i$-th channel as $S_{ii}=\left(\frac{1}{n}\sum_{j=1}^{n}|\tilde{X}_{ij}|\right)^{\frac{1}{2}}$. Here, $n$ denotes the number of activation entries in the $i$-th channel. We then apply SVD to the scaled error $\Delta W^{"}=\Delta WS$ to obtain its low-rank approximation. Since $S$ is invertible, we can project the approximation back to the original space as $\Delta W=\Delta W^{"}S^{-1}\approx B^{"}A^{"}S^{-1}$—same as how EoRA project its compensation back to the original space. We refer to this method as Act-S in the remainder of this paper. We also compare EoRA with a training-based method, ApiQ (liao2024apiq), which optimizes low-rank matrices ($A$ and $B$) using gradient-based training to minimize Eq. 6. In our comparison, we limit the evaluation to the layer-wise variant of ApiQ, as other variants require substantially more memory or training time. These more resource-intensive versions align more closely with PEFT methods rather than fine-tuning-free low-rank approximation approaches. For instance, when applied at the model level, ApiQ effectively becomes equivalent to training LoRA on top of a compressed model—shifting its focus toward fine-tuning rather than fine-tuning-free compensation, and thus falling outside the scope of this study. Note that the optimization time for both EoRA and Act-S is comparable, with each completing within minutes, whereas ApiQ requires over hours to optimize.
@@
We evaluate EoRA and the baselines on improving the task-specific accuracy of the compressed LLMs on language generation, commonsense reasoning, and math reasoning tasks using the LM-Evaluation-Harness framework (eval-harness). We pick WikiText2 for the language generation task and perplexity as the evaluation metric. For commonsense reasoning, we select ARC-Challenge (ARC-C) (clark2018think), and for math reasoning ability, we choose MathQA (amini2019mathqa) and GSM8K (cobbe2021training). We sample 128 concatenated sentences of length 2048 from the WikiText2 training set as the calibration set for EoRA, Act-S, and ApiQ for the language generation task. For commonsense reasoning tasks, we sample 32 concatenated sentences of length 2048 from the ARC training set and combine them with 32 concatenated sentences of the same length from C4 (raffel2020c4) to construct the calibration set for EoRA, Act-S, and ApiQ. Similarly, for the math reasoning tasks, we sample 32 concatenated sentences of length 2048 from the MathQA/GSM8K training set and combine them with 32 concatenated sentences from C4 to form the calibration set for the three methods.
@@ 4.2.1 Sparsity Error Compensation
To assess the effectiveness of EoRA in compensating for sparsity error, we compare EoRA with all the baselines on LLaMA2-7B/13B and LLaMA3-8B models pruned with SparseGPT to 2:4 sparsity—the only sparsity format that yields actual inference speedups on GPUs. Rank of all the compensation methods is set to 128, and the results of LLaMA3-8B are presented in Table 1, while the full results, including LLaMA2-7B/13B, are provided in Table 4 in the appendix. EoRA consistently outperforms all fine-tuning-free baselines, achieving gains of $2.9\%$, $2.1\%$, and $10.7\%$ over Act-S on ARC-C, MathQA, and GSM8K, respectively. Furthermore, it surpasses ApiQ by $0.4\%$ on ARC-C and $1.1\%$ on MathQA, while delivering comparable results on GSM8K—all with significantly faster optimization time (15 minutes vs. 2.5 hours). Furthermore, EoRA proves robustness across different model sizes, continuing to outperform ZeroQuant-V2, Act-S, and ApiQ in boosting the accuracy of 2:4 pruned LLaMA2-7B/13B across ARC-C and MathQA as shown in Table 4. We further assess the generalizability and compatibility of EoRA with pruning methods beyond SparseGPT. Specifically, we evaluate EoRA on LLaMA3-8B pruned to 2:4 sparsity using Wanda (sun2023simple), where EoRA continues to outperform all the fine-tuning-free baseline methods. For additional details, please refer to Section A.4.
@@ 4.2.2 Quantization Error Compensation
We evaluate EoRA on LLaMA2-7B/13B and LLaMA3-8B models quantized with GPTQ to 4-bit and 3-bit to assess the effectiveness of EoRA in compensating for quantization error. The ranks for all the methods are set to 128. From Table 2, 3-bit quantization causes significant accuracy degradation, with losses up to $29.5\%$$17.7\%$$35.8\%$ on ARC-C, MathQA, and GSM8K, respectively. By applying EoRA, we demonstrate that the accuracy loss can be reduced to $18.7\%$$10.9\%$$24.3\%$ on ARC-C, MathQA, and GSM8K—providing $10.8\%$$6.7\%$$11.5\%$ improvement, outperforming all the baseline methods for compensating the quantization error. On the other hand, although 4-bit quantization does not result in as much accuracy loss as 3-bit quantization, applying EoRA can still generally enhance the performance of the 4-bit model, offering up to a $2.2\%$ and $3.14\%$ accuracy boost on ARC-C and MathQA, respectively. Comprehensive results, including those for LLaMA2-7B/13B, are presented in Table 5 in the appendix, where a similar trend of improvement with EoRA is observed. We also explore the feasibility of using EoRA to improve ultra-compressed models that combine both pruning and quantization. In this setting, EoRA continues to outperform all baselines on ARC-C and MathQA. Further details can be found in Section 6 of the Appendix.
@@ 4.3 Ablation Study: Ranks and Calibration Sizes
Since one of the advantages of using EoRA is the greater flexibility in adjusting overall model accuracy without being constrained by specific compression formats, in this section, we investigate the influence of different ranks on adopting EoRA. We vary the rank of EoRA in {64,128,256,512} on compensating LLaMA3-8B pruned to 2:4 sparsity. As shown in Figure 2, EoRA consistently outperforms the two fine-tuning-free baselines (ZeroQuant-V2 and Act-S) across all tested ranks, with the performance gap becoming more prominent at higher ranks. For instance, on GSM8K, EoRA achieves improvements of $7.43\%$, $0.69\%, 11$$1.9\%$, and $14.62\%$ at ranks 64, 128, 256, and 512, respectively. In contrast, the gains on ARC-C remain relatively stable across ranks, ranging between $2\%$ and $4\%$. Additionally, EoRA begins to outperform ApiQ on GSM8K at higher ranks, with improvements of $1.21\%$ and $2.51\%$ observed at ranks 256 and 512, respectively. These experiments prove that EoRA is robust across different rank settings, offering users a more flexible option upon existing compression configurations to effectively balance the trade-off between inference overhead and model accuracy. A similar trend is observed in the results for LLaMA2-7B/13B shown in Table 10 in the appendix.
We also compare the influence of different calibration sizes on EoRA . We vary the calibration size in {16,32,64}, and compare them on recovering the accuracy of LLaMA3-8B quantized to 3/4-bit and pruned to 2:4 sparsity. Overall, we find that EoRA demonstrates strong robustness and maintains competitive accuracy even with limited calibration data, as shown in Table 8. Notably, using as few as 32 calibration samples to compensate for a 2:4 pruned model can even yield better accuracy improvements than using 64 samples.
@@ 4.4 EoRA as LoRA initialization for Fine-tuning Compressed Models
LLaMA3-8B Full-precision - w/o fine-tuning 50.42 40.10
@@
EoRA (Ours) 55.46 56.04
We show that, with additional computational resources, users can leverage the low-rank matrices from EoRA as initialization for LoRA fine-tuning, enabling further accuracy improvements for compressed models. We follow the conventional LoRA fine-tuning framework, which keeps the compressed model frozen and only tunes the low-rank residual components during fine-tuning. We conduct experiments on compressed LLaMA3-8B models with {2:4 sparsity, 4-bit, 3-bit} compression. The rank of LoRA is set to 128 and is applied to every linear layer, initialized using EoRA, SVD following LoftQ (liloftq), and standard initialization following QLoRA (dettmers2024qlora). Fine-tuning is performed on the ARC training set for evaluating ARC-C, and on the MathQA training set for the math reasoning task. We fine-tune the models for 3 epochs with a batch size of 64, a learning rate of 1e-5, and a cosine learning rate scheduler. As shown in Table 3, initializing with EoRA substantially enhances the accuracy of compressed models, surpassing both QLoRA and LoftQ when fine-tuning 4-bit quantized LLaMA3-8B, and achieving accuracy on par with standard full-precision fine-tuning. We also observed that the improvements over QLoRA and LoftQ are more pronounced on 3-bit quantized and 2:4 pruned models, aligning with our earlier finding that EoRA is more effective when the compression error is more substantial, as shown in Table 11.
@@ 4.5 Kernel Optimization, Inference Speed Evaluation and Memory Overhead of EoRA
While theoretically, compensating a compressed model with low-rank residual paths introduces minimal computational overhead, in practice, it leads to a noticeable increase in latency. This is primarily because input and output must transfer between L2 cache and DRAM twice as often compared to that without a low-rank residual path, shifting the inference process from being computation-bound to memory-bound. This phenomenon is also discussed in (li2024svdqunat). To address this, we propose fusing the low-bit weight quantization kernel with the matrix multiplication of $B$, which shares the same output. By doing so, the shared output no longer needs to be offloaded and reloaded to the L2 cache, effectively reducing data transfer overhead as illustrated in Figure 3 (a). Implementation details of our kernel can be found in Section A.6. As shown in Table 9, our custom EoRA kernel substantially accelerates inference compared to using native PyTorch for the low-rank residual path on top of the low-bit quantized kernel, achieving a speedup of up to 1.4x over FP16 with EoRA of rank 128 at 3-bit quantization. In contrast, without the EoRA kernel, the initial 1.7x speedup provided by the 3-bit quantized kernel drops to 1.1x. Similarly, under 4-bit quantization, the EoRA kernel delivers an extra 0.3x speedup compared to setups without the EoRA kernel.
Finally, EoRA can also be quantized to further reduce the additional cost of residual low-rank compensation paths. In this section, we quantize EoRA of rank {128, 512} to 4/3-bit on compensating three types of compressed LLaMA3-8B models (2:4 pruned, 4-bit quantized, and 3-bit quantized). The complete results are provided in Table 13 in appendix, while the results for LLaMA3-8B are illustrated in Figure 3 (b). As shown in the figure, EoRA is robust to quantization, which means that when EoRA is quantized, the accuracy drop from full-precision EoRA is insignificant while the model size is significantly reduced. For example, when a 512-rank EoRA is quantized from 16-bits to 4-bit on 2:4 pruned LLaMA3-8B, the accuracy drops are only $0.43\%$ on ARC-C while the total model size reduces by $16.49\%$. Additionally, compared to the original uncompensated 2:4 pruned model, quantizing EoRA of rank 128/512 improves accuracy by $4.4\%$$11.4\%$ with a total model size increase of just $2\%$$7\%$. For 3-bit quantized LLaMA3-8B compensated with a 4-bit quantized EoRA of rank 128/512 achieves $10.6\%$$19.1\%$ accuracy improvements, with a corresponding model size increase of only $3\%$$14\%$. Interestingly, we also observe that quantizing EoRA does not always result in accuracy loss; in some cases, it even slightly improves accuracy, potentially due to quantization acting as a form of regularization, as discussed in OFQ (liu2023oscillation). Generally, we recommend users quantize EoRA to 4-bit, as this significantly reduces inference latency and model size with kernel support, without causing any noticeable drop in accuracy.
@@ 5 Related Works
Post-training LLM Compression:
As LLMs scale, reducing their size is essential for efficient deployment. Traditional compression-aware training methods are impractical due to the need for full datasets and heavy retraining. Post-training compression methods like quantization and pruning have gained popularity as they require only minimal calibration data and no retraining. PTQ reduces model size by lowering bitwidths (frantar2022gptq; tseng2024quip), while PTP removes less important weights to reduce computation (frantar2023sparsegpt; sun2023simple). Our method, EoRA, is compatible with all such compression techniques as it operates independently of the base method used.
Low-Rank Decomposition:
Low-rank decomposition methods (yuan2023asvd; sakr2024espace; wang2024svdllm; hsu2022fwsvd; mozaffari2024slim; zhang2024oats; zhang2024lqer; zhang2024qera; saha2024compressing) compress models by replacing weights with low-rank matrices, reducing both latency and size without special kernel support. However, they are less widely adopted due to weaker accuracy-compression trade-offs. While FWSVD (hsu2022fwsvd) is designed primarily as a compression method based on low-rank decomposition of model weights, EoRA instead leverages low-rank modules specifically to compensate for compression errors. Unlike FWSVD, EoRA provides a theoretical guarantee of minimizing layer-wise compression loss, as demonstrated in our derivation in Section 3. Using gradient-based information, as required by FWSVD, can be prohibitively expensive for LLMs, as noted in ASVD (yuan2023asvd). SVD-LLM (wang2024svdllm) is conceptually close to our work, which also tries to align the SVD compression error with the layer-wise compression loss. It relies on the matrix product of the activation being positive-definite, a condition often unmet in practice. Enforcing this condition typically requires additional modifications, which introduce noise into the approximation. In contrast, EoRA employs eigendecomposition, which only requires the matrix product of the activation to be symmetric—a property that naturally holds—avoiding such issues. Furthermore, although EoRA also utilizes SVD-based low-rank decomposition, its core objective is fundamentally different. Whereas prior methods aim to replace pre-trained weight matrices with low-rank approximations to reduce model size and inference cost, EoRA focuses on approximating the compression error itself. This allows for improved accuracy recovery in compressed LLMs and provides greater flexibility in balancing accuracy and computational overhead by overcoming the constraints of fixed compression formats. Please refer to Section A.13 in Appendix for more detailed comparison.